OpenStudy (anonymous):

Find x: 2-1/3(3x-9)^3=0

OpenStudy (mertsj):

\[2-\frac{1}{3}(3x-9)^3=0\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

1. Subtract 2 from both sides. 2. Multiply both sides by -3 3. Take the third root of both sides 4. Add 9 to both sides 5. Divide both sides by 3

OpenStudy (anonymous):

I got 4.26 as an answer but in the textbook it has 3.6

OpenStudy (mertsj):

\[\frac{-1}{3}(3x-9)^3=-2\] \[(3x-9)^3=8\] \[3x-9=2\] \[3x=11\] \[x=\frac{11}{3}=3 \frac{2}{3}=3.666666666....\]

OpenStudy (anonymous):

Where did you get the 8 from?

OpenStudy (anonymous):

What i did was multiplying the -2 with 3

OpenStudy (mertsj):

yep. I made a mistake. It should be 6

OpenStudy (mertsj):

\[(3x-9)^3=6\] \[3x-9=\sqrt[3]{6}\] \[x-3=\frac{\sqrt[3]{6}}{3}\] \[x=3+\frac{\sqrt[3]{6}}{3}\]

OpenStudy (anonymous):

I did it a different way

OpenStudy (mertsj):

I only did that because I don't have my calculator with me.

OpenStudy (anonymous):

-1/3(3x-9)^3 = -2 -(3x-9)^3 = -6 -3(x-3)^3 = -6 (x-3)^3 = 2

OpenStudy (mertsj):

you cannot factor that 3 out.

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Since it is raised to a power?

OpenStudy (mertsj):

yes